# Proven Approximation Guarantees in Multi-Objective Optimization: SPEA2 Beats NSGA-II

### v1

@@ Abstract
Keywords: Evolutionary algorithms, runtime analysis, SPEA2, OMM, approximation.
@@ 1 Introduction
Many optimization problems in practice consist of several, usually conflicting objectives. One common solution concept for such problems is to compute a set of solutions witnessing the Pareto front (or a sufficiently diverse subset thereof) and then let a human decision make select the final solution. For such multi-objective optimization problems, evolutionary algorithms with their population-based nature are an obvious choice, and in fact, such multi-objective evolutionary algorithms (MOEAs) are among the most successful algorithms for many multi-objective problems Coello et al. (2007); Zhou et al. (2011).
@@
As for other topics in heuristics, also for MOEAs the majority of the research is empirical. However, theoretical analyses of MOEAs, usually proving runtime guarantees and from this aiming at a deeper understanding of these algorithms, exist for more than twenty years now Laumanns et al. (2002); Giel (2003); Thierens (2003). This field has made a huge step forward recently with the first mathematical runtime analysis of the NSGA-II Zheng et al. (2022), the most prominent MOEA. This work was quickly followed up by analyses of other prominent MOEAs such as the NSGA-III Wietheger and Doerr (2023), SMS-EMOA Bian et al. (2023), and SPEA2 Ren et al. (2024).
@@
Interestingly, these and the subsequent results show very similar performance guarantees for these algorithms (which agree with the results known for the classic (G)SEMO analyzed in the early theoretical works on MOEAs). There is one exception, and this is the result Zheng and Doerr (2024b) showing that the NSGA-II has enormous difficulties for three or more objectives. This problem of the NSGA-II, however, can easily be resolved with an additional tie-breaker and then again the same performance guarantees as known for the other algorithms hold.
@@
In this work, for the first time among these common MOEAs, we detect a notable performance difference, namely in the ability to approximate the Pareto front. This aspect is understood much less than the time to compute the full Pareto front. The only such result for the algorithms mentioned above is Zheng and Doerr (2024a). In that work, it was argued via theoretical arguments that the classic NSGA-II, computing the crowding distance for all individuals and then selecting the next population, can have difficulties to approximate the Pareto front as it ignores that every removal of an individual affects the crowding distance of other solutions. However, for the algorithm variant that removes the individuals sequentially and updates the crowding distances after each removal (as suggested already in the little known work Kukkonen and Deb (2006)), a $2$-approximation result was shown for the optimization of the bi-objective OneMinMax benchmark. This is the only approximation guarantee so far for one of the classic algorithms named above.
@@ 2 Preliminaries
We denote the natural numbers, including 00, by $\mathbb{N}$ and the real numbers by $\mathbb{R}$. For all $m,n\in\mathbb{N}$, we let $[m..n]\coloneqq[m,n]\cap\mathbb{N}$ and $[n]\coloneqq[1..n]$.
For all $n\in\mathbb{N}_{\geq 1}$, we call $x\in\{0,1\}^{n}$ an individual, denoting each position $i\in[n]$ by $x_{i}$. We call a multi-set of individuals a population, and we use standard set operations for populations, which extend naturally to multi-sets. We consider pseudo-Boolean maximization of bi-objective functions $f\colon\{0,1\}^{n}\to\mathbb{R}^{2}$, which map individuals to objective values. For each $x\in\{0,1\}^{n}$ and $i\in[2]$, we denote the objective value $i$ of $x$ by $f_{i}(x)$.
We compare objective values $u,v\in\mathbb{R}$ via the dominance partial order. We say that $u$ weakly dominates $v$ (written $u\succeq v$) if and only if $u_{1}\geq v_{1}$ and $u_{2}\geq v_{2}$, and we say that $u$ strictly dominates $v$ if and only if at least one of these inequalities is strict. If and only if neither $u\succeq v$ nor $v\succeq u$, we say that $u$ and $v$ are incomparable. Given a bi-objective function, we extend this notion to all individuals by implicitly referring to their objective value.
For a (multi-)set $P$ of individuals and a bi-objective function $f$, we say that an individual $x\in P$ is non-dominated if and only if there is no $y\in P$ that strictly dominates $u$. We call non-dominated individual of $\{0,1\}^{n}$ Pareto-optimal and its objective value a Pareto optimum. Last, we call the set of all Pareto optima the Pareto front (of $f$).
We use traditional set notation for both normal sets (with unique elements) and multi-sets (that may contain duplicates). The union of multi-sets does not remove any duplicates, and the cardinality of a multi-set accounts for all duplicates. However, if we apply a function to multi-set, then the result is a normal function, that is, the function values are not duplicated. Note that populations are multi-sets whereas sets of objective values are normal sets.
2.1 The OneMinMax Benchmark
@@
such that $\textrm{such that }\alpha\beta+(\alpha+1)(\mu-1-\beta)=n,$
@@ 3.1 The SPEA2
The SPEA2 maximizes a given bi-objective problem $f\colon\{0,1\}^{n}\to\mathbb{R}$ by iteratively refining a multi-set of individuals (the parent population) of given population size $\mu\in\mathbb{N}_{\geq 1}$. In each iteration, the algorithm generates a user-defined amount $\lambda\in\mathbb{N}_{\geq 1}$ of new individuals (the offspring population) via a process called mutation. Afterward, the SPEA2 selects the $\mu$ most promising individuals among the parent and the offspring population, to be used in the next iteration. The basis for this new population are all non-dominated individuals. If this number is greater than $\mu$, then the algorithm removes individuals based on a clustering technique called $\sigma$-distances. This is the predominant case in our analysis in Section 4, as the OMM benchmark only features non-dominated objective values.
@@
Given a population $P$ of size $s\in\mathbb{N}_{\geq 1}$ of non-dominated individuals, the $\sigma$-distances are based on a function $\sigma\colon\{0,1\}^{n}\to\mathbb{R}^{s-1}$ that assigns each individual $x\in P$ a vector that contains the Euclidean distances of $x$ to all other individuals in $P$ in ascending order (breaking ties arbitrarily), which we call the $\sigma$-distance of $x$. We follow the convention of Zitzler et al. (2001) and denote for all $i\in[s-1]$ the distance of $x$ to its $i$-closest neighbor by $\sigma^{i}_{x}$.
@@ 1 $t\leftarrow 0$;
4       $Q_{t}\leftarrow$ offspring population of $P_{t}$ of size $N$;
@@ 3.2 The NSGA-II
Like the SPEA2, the NSGA-II maintains a parent population of $N\in\mathbb{N}_{\geq 1}$ individuals, from which it generates $N$ offspring each iteration, using a mutation operator. Afterward, it selects individuals greedily based on the number of individuals by which they are strictly dominated, starting with the non-dominated individuals. During this phase, all individuals with an identical number are selected. The smallest number where this leads to selecting at least $N$ individuals in total is known as the critical rank. For OMM, this means that the entire combined parent and offspring population (of size $2N$) is kept, as all individuals are not strictly dominated, thus all land in the critical rank.
@@ 4 The SPEA2 Computes an Optimal Spread Efficiently
Our main result is Theorem 1, which shows that the steady-state SPEA2 with 1-bit mutation and $\mu\leq\frac{n}{3}$ computes an optimal spread on OMM after $O(\mu^{2}n\log(\mu)\log(n))$ expected function evaluations [^1] This is asmyptotically equal to the number of iterations, as each iteration generates exactly one offspring in the steady-state variant.. This bound is slower by a factor of $\mu\log\mu$ than the time required to find the extreme objective values (Theorem 4), which may be an artifact of our analysis.
@@
In the following, we start with a high-level overview of the proof structure. Afterward, we provide more details, until the prove Theorem 1.
Due to space limitations, most of our proofs can be found in the technical appendix of this paper.
@@ 4.1 High-Level Proof Outline for the SPEA2
In order to remove a minimum length interval, it is sufficient to have it neighbor an interval whose length is at least two larger. By placing an offspring in between these two intervals such that it reduces the size of the larger interval, the size of the minimum length interval increases by one if the parent is removed. Lemma 10 bounds the expected number of iterations it takes to remove one minimum length interval like this. In a nutshell, the analysis relies on the fact that minimum length intervals tend to move to the borders ot the objective space $[0..n]$. Once there, the minimum distance between a minimum length interval and one with a length at least two larger does not increase. This allows eventually to decrease the number of intervals of minimum length.
@@ 4.2 The SPEA2 Computes an Optimal Spread Efficiently
For 1-bit mutation and $\mu\leq\frac{n}{3}$, this is asymtpotically tight.
@@
Due to space restrictions, we detail the underlying case distinction in the supplementary material. Our main result following from the discussion above is the following bound on the number of iterations to increase the minimum interval length or the number of intervals thereof.
@@
We combine all of our prior arguments in order to prove Theorem 1. We essentially rely on Lemma 10 to gradually increase the minimum interval length until it reaches the optimal value $\alpha$ (equation 1). However, this requires that the minimum interval length is at least $2$. By Lemma 9, we achieve such a state quickly, but it requires us to restrict the population size $\mu$ of the SPEA2 to $\frac{n}{3}$.
@@ 5 The NSGA-II Does Not Compute an Optimal Spread in Polynomial Time
We show that the steady-state NSGA-II takes with high probability a super-polynomial time to compute an optimal spread when starting in an unfavorable state. This is in stark contrast to the SPEA2, which always computes an optimal spread in a rather fast polynomial time (Theorem 1).
@@
Note that the only possible changes to the population are those that change where the intervals of length $c+1$ or $c-1$ are located as well as the change that averages the two intervals of lengths $c+1$ and $c-1$ to $c$. In the latter case, the algorithm computes an optimal spread. However, to do so, it is necessary that those two intervals are next to each other. Before this is the case, they can only move their interval index by one by swapping their position with a neighboring interval of length $c$.
Let $T$ be the first iteration such that the two intervals of lengths $c+1$ and $c-1$ are neighbors. We show that the probability that $T$ is polynomial is at most $e^{-\Omega(n)}$, thus proving the claim. Hence, in the following, we always assume implicitly that we only consider iterations $t\in T$ such that $t<T$ [^2] Formally, this requires multiplying each statement with the indicator variable of the event $\{t<T\}$, which we omit.
For all $t\in\mathbb{N}$, let $X_{t}$ denote the index of the interval of length $c+1$ at the beginning of iteration $t$. (Note that we omit the case if no such interval exists, as we only consider iterations less than $T$.) Analogously, let $Y_{t}$ denote the index of the interval of length $c-1$ at the beginning of iteration $t$. Note that both $X$ and $Y$ change by at most $1$ in each iteration.
We first consider $X$ and determine its expected change, aiming to apply Theorem 12. To this end, we only consider those iterations in which $X$ actually changes (and iteration 00) and only while $X$ is in $[\frac{1}{16}n^{\prime}..\frac{3}{16}n^{\prime}]$. Moreover, we assume that the interval of length $c-1$ does not move. Let the resulting process be $X^{\prime}$, and let $T_{x}$ be the first point in time $t\in\mathbb{N}$ of process $X^{\prime}$ such that $X^{\prime}\geq\frac{3}{16}n^{\prime}$. Note that $X^{\prime}_{0}=\frac{2}{16}n^{\prime}$ by assumption. Furthermore note that for all $t\in\mathbb{N}$, if $X^{\prime}_{t}<\frac{3}{16}n^{\prime}$, the same is true for $X$. Hence, if $Y\geq\frac{3}{16}n^{\prime}$ until at least $T_{x}$, then not considering the interval of length $c-1$ moving does not affect the statements about $X^{\prime}$ at all. We make use of this observation at the end of this proof.
@@
For $Y$ we proceed analogously, defining $Y^{\prime}$ in the same way as $X^{\prime}$ but only consider iterations where $Y$ is in $[\frac{3}{16}n^{\prime}..\frac{5}{16}n^{\prime}]$. Moreover, let $T_{y}$ be the first point in time $t\in\mathbb{N}$ of process $Y^{\prime}$ such that $Y^{\prime}_{t}\leq\frac{3}{16}n^{\prime}$. Last, let $t\in\mathbb{N}$ (less than $T_{y}$) and let $d^{\prime}\coloneqq\bigl{(}(n-cY^{\prime}_{t})/n\bigr{)}\cdot\frac{1}{2}+\bigl{(}(c(Y^{\prime}_{t}-1)+1)/n\bigr{)}\cdot\frac{1}{2}\leq\frac{1}{2}$. For similar reasons as before, we get
@@
By Theorem 12 applied to $(\frac{4}{16}n^{\prime}-Y^{\prime}_{t})_{t\in\mathbb{N}}$, we get for all $t\in\mathbb{N}$ that
$\Pr[T_{y}\leq t]\leq t^{2}\exp\Bigl{(}-\frac{(1/16)n^{\prime}(1/2)}{2}\Bigr{)}=t^{2}e^{-\Omega(n)}.$
Overall, we see that $X^{\prime}\geq\frac{3}{16}n^{\prime}$ and $Y^{\prime}\leq\frac{3}{16}n^{\prime}$, which is a necessary condition for the steady-state NSGA-II to compute an optimal spread, occurs for any polynomial number of iterations only with probability $e^{-\Omega(n)}$. Note that since we consider the intersection ot the events that $X^{\prime}<\frac{3}{16}n^{\prime}$ and $Y^{\prime}>\frac{3}{16}n^{\prime}$ for all iterations less than, respectively, $T_{x}$ and $T_{y}$, it does not matter that we did not consider the other interval of length not equal to $c$ for the analysis of either process. Ultimately, since the original process accounts for even more iterations than those considered by $X^{\prime}$ and $Y^{\prime}$, the result follows. ∎

### v3

@@ 1 Introduction
Many optimization problems in practice consist of several conflicting objectives. One common approach for such problems is to compute a set of solutions witnessing the Pareto front (or a sufficiently diverse subset thereof) and then let a human decision maker select the final solution. For such multi-objective optimization problems, evolutionary algorithms with their population-based nature are an obvious choice, and in fact, such multi-objective evolutionary algorithms (MOEAs) are among the most successful algorithms Coello et al. (2007); Zhou et al. (2011).
@@
Most research on randomized search heuristics is empirical, and much fewer theoretical works exist Neumann and Witt (2010); Auger and Doerr (2011); Jansen (2013); Zhou et al. (2019); Doerr and Neumann (2020). However, theoretical analyses of MOEAs, usually proving runtime guarantees and from this aiming at a deeper understanding of these algorithms, exist for more than twenty years Laumanns et al. (2002); Giel (2003); Thierens (2003). This field has made a huge step forward recently with the first mathematical runtime analysis of the NSGA-II Zheng et al. (2022); Zheng and Doerr (2023), the most prominent MOEA. This work was quickly followed up by more detailed analyses of the NSGA-II Bian and Qian (2022); Doerr and Qu (2023a, b, c); Dang et al. (2023); Cerf et al. (2023); Dang et al. (2024); Deng et al. (2024); Zheng and Doerr (2024a, b); Doerr et al. (2025) and by analyses of other prominent MOEAs… such as the NSGA-III Wietheger and Doerr (2023); Opris et al. (2024); Deng et al. (2025); Opris (2025), SMS-EMOA Bian et al. (2023); Zheng and Doerr (2024c); Li et al. (2025), and SPEA2 Ren et al. (2024).
@@
Interestingly, these results show very similar performance guarantees for these algorithms (which agree with the results known for the classic (G)SEMO analyzed in the early theoretical works on MOEAs). Also, Ren et al. (2024) provide a mathematical framework allowing to prove comparable runtime bounds for several MOEAs. The sole outlier so far is the result Zheng and Doerr (2024b) showing that the NSGA-II has enormous difficulties for three or more objectives (see Doerr et al. (2024) for a second such result). This problem of the NSGA-II can be resolved with an additional tie-breaker Doerr et al. (2025) and then the same performance guarantees as known for the other algorithms hold.
@@
In this work, we detect a second notable performance difference, namely in the ability to approximate the Pareto front. This aspect is understood much less than the time to compute the full Pareto front. The only such result for the algorithms mentioned above is Zheng and Doerr (2024a). In that work, it was argued via theoretical arguments that the classic NSGA-II, computing the crowding distance for all individuals and then selecting the next population, can have difficulties to approximate the Pareto front as it ignores that every removal of an individual affects the crowding distance of other solutions. However, for the algorithm variant that removes the individuals sequentially and updates the crowding distances after each removal (as suggested already in the little known work Kukkonen and Deb (2006)), a $2$-approximation result was shown for the optimization of the bi-objective OneMinMax benchmark. This is the only approximation guarantee so far for one of the classic algorithms named above.
@@ 2 Preliminaries
Denote the natural numbers by $\mathbb{N}$ (with 0) and the reals by $\mathbb{R}$. For all $m,n\in\mathbb{N}$, let $m..n]\coloneqq[m,n]\cap\mathbb{N}$ and $n]\coloneqq[1..n]$.
For all $n\in\mathbb{N}_{\geq 1}$, we call $x\in\{0,1\}^{n}$ an individual. We call a multi-set of individuals a population, and we use standard set operations for populations, which extend naturally to multi-sets. We consider pseudo-Boolean maximization of bi-objective functions $f\colon\{0,1\}^{n}\to\mathbb{R}^{2}$, which map individuals to objective values. For each $x\in\{0,1\}^{n}$ and $i\in[2]$, we denote the objective value $i$ of $x$ by $f_{i}(x)$.
We compare objective values $u,v\in\mathbb{R}$ via the dominance partial order. We say $u$ weakly dominates $v$ (written $u\succeq v$) if and only if $u_{1}\geq v_{1}$ and $u_{2}\geq v_{2}$, and $u$ strictly dominates $v$ if and only if at least one of these inequalities is strict. If and only if neither $u\succeq v$ nor $v\succeq u$, we say $u$ and $v$ are incomparable. Given a bi-objective function, we extend this notion to all individuals by implicitly referring to their objective value.
For a (multi-)set $P$ of individuals and a bi-objective function $f$, we say that an individual $x\in P$ is non-dominated if and only if there is no $y\in P$ that strictly dominates $u$. We call each non-dominated individual of $\{0,1\}^{n}$ Pareto-optimal and its objective value a Pareto optimum. Last, we call the set of all Pareto optima the Pareto front (of $f$).
We use traditional set notation for both normal sets (with unique elements) and multi-sets. The union of multi-sets does not remove any duplicates, and the cardinality of a multi-set accounts for all duplicates. However, if we apply a function to multi-set, then the result is a normal function, that is, the function values are not duplicated. Note that populations are multi-sets whereas sets of objective values are normal sets.
The OneMinMax Benchmark.
@@
$\textrm{such that }\alpha\beta+(\alpha+1)(\mu-1-\beta)=n,$
@@ 3.1 The SPEA2
We regard an equivalent formulation of the SPEA2 that was proposed in the extension of Wietheger and Doerr (2024) (currently only on arXiv). The SPEA2 (Algorithm 1) maximizes a given bi-objective problem $f\colon\{0,1\}^{n}\to\mathbb{R}^{2}$ by iteratively refining a multi-set of individuals (the parent population) of given population size $\mu\in\mathbb{N}_{\geq 1}$. In each iteration, the algorithm generates a user-defined amount $\lambda\in\mathbb{N}_{\geq 1}$ of new individuals (the offspring population) via a process called mutation. Afterward, the SPEA2 selects the $\mu$ most promising individuals among the parent and the offspring population, to be used in the next iteration. The basis for this new population are all non-dominated individuals. If this number is greater than $\mu$, then the algorithm removes individuals based on a clustering technique called $\sigma$-distances. This is the predominant case in our analysis in Section 4, as the OMM benchmark only features non-dominated objective values.
@@
Given a population $P$ of size $s\in\mathbb{N}_{\geq 1}$ of non-dominated individuals, the $\sigma$-distances are based on a function $\sigma\colon\{0,1\}^{n}\to\mathbb{R}^{s-1}$ that assigns each individual $x\in P$ a vector that contains the Euclidean distances of the objective value of $x$ to those of all other individuals in $P$ in ascending order (breaking ties arbitrarily), which we call the $\sigma$-distance of $x$. We follow the convention of Zitzler et al. (2001) and denote for all $i\in[s-1]$ the distance of $x$ to its $i$-closest neighbor by $\sigma^{i}_{x}$.
@@ 1 $t\leftarrow 0$;
4 $Q_{t}\leftarrow$ offspring population of $P_{t}$ of size NN;
@@ 3.2 The NSGA-II
Like the SPEA2, the NSGA-II (Algorithm 2) maximizes a given bi-objective function $f\colon\{0,1\}^{n}\to\mathbb{R}^{2}$ by maintaining a parent population of $N\in\mathbb{N}_{\geq 1}$ individuals, from which it generates $N$ offspring each iteration, using a mutation operator. Afterward, it selects individuals greedily based on the number of individuals by which they are strictly dominated, starting with the non-dominated individuals. During this phase, all individuals with an identical number are selected. The smallest number where this leads to selecting at least $N$ individuals in total is known as the critical rank. For OMM, this means that the entire combined parent and offspring population (of size $2N$) is kept, as all individuals are not strictly dominated, thus all land in the critical rank.
@@ 4 The SPEA2 Computes an Optimal Spread Efficiently
Our main result is Theorem 1, which shows that the steady-state SPEA2 with 1-bit mutation and $\mu\leq\frac{n}{3}$ computes an optimal spread on OMM after $O(\mu^{2}n\log(\mu)\log(n))$ expected function evaluations [^1] This is asymptotically equal to the number of iterations, as each iteration generates exactly one offspring in the steady-state variant.. This bound is slower by a factor of $\mu\log\mu$ than the time required to find the extreme objective values (Theorem 4), which may be an artifact of our analysis.
@@
In the following, we start with a high-level overview of the proof structure. Afterward, we provide more details, until we prove Theorem 1.
@@ 4.1 High-Level Proof Outline for the SPEA2
In order to remove a minimum length interval, it is sufficient to have it neighbor an interval whose length is at least two larger. By placing an offspring in between these two intervals such that it reduces the size of the larger interval, the size of the minimum length interval increases by one if the parent is removed. Lemma 10 bounds the expected number of iterations it takes to remove one minimum length interval like this. In a nutshell, the analysis relies on the fact that minimum length intervals tend to move to the borders of the objective space $[0..n]$. Once there, the minimum distance between a minimum length interval and one with a length at least two larger does not increase. This allows eventually to decrease the number of intervals of minimum length.
@@ 4.2 The SPEA2 Computes an Optimal Spread Efficiently
For 1-bit mutation and $\mu\leq\frac{n}{3}$, this is asymptotically tight.
@@
Our main result following from the discussion above is the following bound on the number of iterations to increase the minimum interval length or the number of intervals thereof.
@@
We combine all of our prior arguments in order to prove Theorem 1. We essentially rely on Lemma 10 to gradually increase the minimum interval length until it reaches the optimal value $\alpha$ (equation 1). However, this requires that the minimum interval length is at least $2$. By Lemma 9, we achieve such a state quickly, but it requires us to restrict the population size $\mu$ of the SPEA2 to at most $\frac{n}{3}$.
@@ 5 The NSGA-II Does Not Compute an Optimal Spread in Polynomial Time
We show that the steady-state NSGA-II takes with high probability a super-polynomial time to compute an optimal spread when starting in an unfavorable state. This is in stark contrast to the SPEA2, which always computes an optimal spread in a rather fast polynomial time (Theorem 1). We note that situations similar to the one we discuss seem to be common, as evidenced in (Zheng and Doerr, 2024a, Figure 1).
@@
Note that the only possible changes to the population are those that change where the intervals of length $c+1$ or c1c-1 are located as well as the change that averages the two intervals of lengths $c+1$ and c1c-1 to $c$. In the latter case, the algorithm computes an optimal spread. However, to do so, it is necessary that those two intervals are next to each other. Before this is the case, they can only move their interval index by one by swapping their position with a neighboring interval of length $c$.
Let $T$ be the first iteration such that the two intervals of lengths $c+1$ and c1c-1 are neighbors. We show that the probability that $T$ is polynomial is at most $e^{-\Omega(n)}$, thus proving the claim. Hence, in the following, we always assume implicitly that we only consider iterations $t\in T$ such that t<Tt<T. [^2] Formally, this requires multiplying each statement with the indicator variable of the event $\{t<T\}$, which we omit.
For all $t\in\mathbb{N}$, let $X_{t}$ denote the index of the interval of length $c+1$ at the beginning of iteration $t$. (Note that we omit the case if no such interval exists, as we only consider iterations less than $T$.) Analogously, let $Y_{t}$ denote the index of the interval of length c1c-1 at the beginning of iteration $t$. Note that both $X$ and $Y$ change by at most $1$ in each iteration.
We first consider $X$ and determine its expected change, aiming to apply Theorem 12. To this end, we only consider those iterations in which $X$ actually changes (and iteration 0) and only while $X$ is in $\frac{1}{16}n^{\prime}..\frac{3}{16}n^{\prime}]$. Moreover, we assume that the interval of length c1c-1 does not move. Let the resulting process be $X^{\prime}$, and let $T_{x}$ be the first point in time $t\in\mathbb{N}$ of process $X^{\prime}$ such that $X^{\prime}\geq\frac{3}{16}n^{\prime}$. Note that $X^{\prime}_{0}=\frac{2}{16}n^{\prime}$ by assumption. Furthermore note that for all $t\in\mathbb{N}$, if $X^{\prime}_{t}<\frac{3}{16}n^{\prime}$, the same is true for $X$. Hence, if $Y\geq\frac{3}{16}n^{\prime}$ until at least $T_{x}$, then not considering the interval of length c1c-1 moving does not affect the statements about $X^{\prime}$ at all. We make use of this observation at the end of this proof.
@@
For $Y$ we proceed analogously, defining $Y^{\prime}$ in the same way as $X^{\prime}$ but only consider iterations where $Y$ is in $\frac{3}{16}n^{\prime}..\frac{5}{16}n^{\prime}]$. Moreover, let $T_{y}$ be the first point in time $t\in\mathbb{N}$ of process $Y^{\prime}$ such that $Y^{\prime}_{t}\leq\frac{3}{16}n^{\prime}$. Last, let $t\in\mathbb{N}$ (less than $T_{y}$) and let $d^{\prime}\coloneqq\bigl{(}(n-cY^{\prime}_{t})/n\bigr{)}\cdot\frac{1}{2}+\bigl{(}(c(Y^{\prime}_{t}-1)+1)/n\bigr{)}\cdot\frac{1}{2}\leq\frac{1}{2}$. For similar reasons as before, we get
@@
By Theorem 12 applied to $(\frac{4}{16}n^{\prime}-Y^{\prime}_{t})_{t\in\mathbb{N}}$, we get for all $t\in\mathbb{N}$ that $\Pr[T_{y}\leq t]\leq t^{2}\exp(-\frac{(1/16)n^{\prime}(1/2)}{2})=t^{2}e^{-\Omega(n)}$.
Overall, we see that $X^{\prime}\geq\frac{3}{16}n^{\prime}$ and $Y^{\prime}\leq\frac{3}{16}n^{\prime}$, which is a necessary condition for the steady-state NSGA-II to compute an optimal spread, occurs for any polynomial number of iterations only with probability $e^{-\Omega(n)}$. Note that since we consider the intersection of the events that $X^{\prime}<\frac{3}{16}n^{\prime}$ and $Y^{\prime}>\frac{3}{16}n^{\prime}$ for all iterations less than, respectively, $T_{x}$ and $T_{y}$, it does not matter that we did not consider the other interval of length not equal to $c$ for the analysis of either process. Ultimately, since the original process accounts for even more iterations than those considered by $X^{\prime}$ and $Y^{\prime}$, the result follows. ∎